OpenStudy (anonymous):

Which of the following equations is parallel to the line passing through A(−2, 7), B(−2, −2)?  (1 point) y = −2 y = x 3x = 5 −2x = 7y

OpenStudy (anonymous):

Same type of question as before...let me see your work.

OpenStudy (shamil98):

A(−2, 7), B(−2, −2) first find the slope remember parallel lines have the same slope :)

OpenStudy (anonymous):

its -2.. right?

OpenStudy (anonymous):

NO, it is (-2-7)/(-2 - -2) = -9/0 which is undefined slope

OpenStudy (shamil98):

A(−2, 7), B(−2, −2)? (1 point) -2 -(-2) / -2 - 7 = -2 + 2 / -2- 7 = 0 / -9 = 0

OpenStudy (shamil98):

It is not undefined @Easyaspi314 ...

OpenStudy (anonymous):

Shamil...your definition of slope is INCORRECT. Look at how I found the slope!

OpenStudy (shamil98):

y2 - y1/ x2 -x1

OpenStudy (shamil98):

How am i the incorrect one?

OpenStudy (anonymous):

So that gives youa 0 in the denominator. Slope is UNDEFINED!

OpenStudy (anonymous):

x2 - x1 = 0

OpenStudy (anonymous):

Well????????? See your error?